Assume that Ocean King Products sells three varieties of canned seafood with the following prices and costs.

Selling Price per case Variable Cost per Case Fixed Cost per Month

Variety 1 $ 23 $ 19 –

Variety 2 24 22 –

Variety 3 29 20 –

Entire firm – – $ 50,000

The sales mix (in cases) is 50%  Variety 1, 25% Variety 2, and 25% Variety 3.

a. At what sales revenue per month does the company break even?

b. Suppose the company is subject to a 35 percent tax rate on income. At what sales revenue per month will the company earn $57,005 after taxes assuming the same sales mix?
Answer #1

Ocean King Products

Weighted average contribution per unit: ($)

Variety Selling price (S) Variable cost (V) Contribution per unit (C = S - V) Sales mix (M) (C×M)
1 23 19 4 50% 2
2 24 22 2 25% 0.5
3 29 20 9 25% 2.25
$ 4.75

a) Break even sales per month:

Break even point = Fixed cost / Weighted average contribution per unit

= $ 50,000/$ 4.75 = 10,526 units

Variety BEP units × Sales mix × Selling price $
1 10,526 × 50% × $ 23 121,049
2 10,526 × 25% × $ 24 63,174
3 10,526 × 25% × $ 29 76,335
$ 260,558

b) Sales revenue per month (target income $ 57,005 after tax = $ 87,700 before tax (57,005 ÷ (1-0.35)))

* Targeted income per month $ 7,308.33, assuming $ 87,700 is for 12 months

= (Fixed cost + Targeted income) / weighted average contribution per unit

= (50,000+7,308.33)/4.75 = 12,065 units

Variety BEP units × Sales Mix × Selling price $
1 12065 × 50% × $ 23 138,748
2 12065 × 25% × $ 24 72,390
3 12065 × 25% × $ 29 87,471
$ 298,609